OpenStudy (anonymous):

Using integrals to find the average value of a function over a given interval.

OpenStudy (anonymous):

I think the answer to this question is 0.

OpenStudy (anonymous):

\[\frac{1}{b-2}\int_2^bf(x)dx=4\] and \[\int_2^bf(x)dx=8\]

OpenStudy (anonymous):

so \[b-2=2\implies b=4\] yup looks good

OpenStudy (anonymous):

oh you wrote 0 i think it is 4 check should work

OpenStudy (anonymous):

oh....right! I mistook -1/2 for 1/2. It would be 4. Thanks satellite73, you rock!

OpenStudy (anonymous):

yw